OpenStudy (anonymous):

1.A 6 month loan with interest of $157 is paid in full with 1 payments remaining. How musch interest should be refunded to the borrower?

OpenStudy (anonymous):

157/6 = 26.17*1 = 26.17 for the one month remaining.

OpenStudy (anonymous):

thank you